Give the functions for the following relations: (1,1) , (3,11) , (5,21) , (7,31) , (9,41).
Answer by stanbon(75887)

Give the functions for the following relations: (1,1) , (3,11) , (5,21) , (7,31) , (9,41).
-------
As x increases 2, y increases 10
----
slope = 10/2 = 5
----
Form: y = mx + b
Using the point (1,1)
1 = 5*1 + b
b = -4
---
Equation:
y = 5x-4
==============
Cheers,
Stan H.
==============